# Constraining interacting dark energy models with black hole superradiance

### v1

@@ 5.2 DE superradiance induced by DM spike
$\Omega(\Phi)=1+\frac{1}{2}\beta\left(\frac{\Phi}{M_{\mathrm{Pl}}}\right)^{2}.$ (5.2)
@@
$\square_{\mathrm{Kerr}}\Phi=\mu^{2}_{\mathrm{eff}}(\mathbf{r})\Phi,$ (5.3)
@@
$\mu_{\mathrm{eff}}^{2}(\mathbf{r})=\mu_{0}^{2}+\frac{\beta}{M_{\mathrm{Pl}}^{2}}\rho_{\mathrm{DM}}(\mathbf{r})\,.$ (5.4)
@@
However, the immense gravitational pull of an SMBH can dramatically increase the DM density in its vicinity, forming a structure known as a DM spike, where $\rho_{\mathrm{spike}}\gg\rho_{\mathrm{local}}$ [PhysRevLett.83.1719]. This extreme density enhancement provides a mechanism to significantly boost the DE effective mass. A DM halo with an initial central density profile $\rho(r)=\rho_{0}(r/r_{0})^{-\gamma}$ will form a spike around an SMBH with a profile given by [PhysRevLett.83.1719]
$\rho_{\mathrm{sp}}(r)=\rho_{R}\Theta(r-4R_{s})\left(1-\frac{4R_{\mathrm{s}}}{r}\right)^{3}\left(\frac{R_{\mathrm{sp}}}{r}\right)^{\gamma_{\mathrm{sp}}},$ (5.5)
@@
where $R_{\mathrm{s}}=2GM$ is the Schwarzschild radius of the SMBH; $\gamma_{\mathrm{sp}}=(9-2\gamma)/(4-\gamma)$ which yields $2.25<\gamma_{\mathrm{sp}}<2.5$ for $0<\gamma<2$. $\rho_{R}$ and $R_{\mathrm{sp}}$ is the spike parameter which depend on $\gamma$ and the BH mass $M$ through the relation $R_{\mathrm{sp}}\left(\gamma,M\right)=\alpha_{\gamma}r_{0}\left(M/\left(\rho_{0}r_{0}^{3}\right)\right)^{1/(3-\gamma)}$ and $\rho_{R}=\rho_{0}\left(R_{\mathrm{sp}}/r_{0}\right)^{-\gamma}$, where the normalization $\alpha_{\gamma}$ is numerically derived for each power-law index $\gamma$, e.g., $\alpha_{1}\approx 0.1$ and $\alpha_{2}\approx 0.02$ [PhysRevLett.83.1719]. We determine the initial $\rho_{0}$ and $r_{0}$ through the mass-velocity-dispersion relationship following [Nishikawa:2017chy] (see Appendix C for more details).
@@
Figure 2: Dark matter density profiles around a SMBH with the mass of M87* ($M=6.5\times 10^{9}\mathrm{M}_{\odot}$). The solid line represents the DM spike profile resulting from an initial NFW halo ($\gamma=1$, $\gamma_{\mathrm{sp}}=7/3$), while the dashed line shows the steeper spike resulting from an initial $\gamma=2$ halo ($\gamma_{\mathrm{sp}}=5/2$). The density is shown as a function of the radial coordinate $r$. The density is shown as a function of the radial coordinate $r$, covering the range from the event horizon $r_{+}$ to the spike radius $R_{\mathrm{sp}}$.
@@
In this work, we will consider two representative cases for the initial halo profile. The first is the standard Navarro-Frenk-White (NFW) profile [Dubinski:1991bm], corresponding to $\gamma=1$, which results in a spike density that scales as $\rho_{\mathrm{sp}}(r)\propto r^{-7/3}$. As a second, more optimistic case for generating a strong superradiance signal, we consider a steeper initial profile with $\gamma=2$, leading to an even denser spike with $\rho_{\mathrm{sp}}(r)\propto r^{-5/2}$. Since a steeper initial profile produces a denser DM spike, a smaller value of the coupling $\beta$ is required to enhance the DE effective mass to the superradiance threshold. We will derive constraints for both the $\gamma=1$ and $\gamma=2$ scenarios. The resulting DM spike density profiles for these two cases, calculated using the mass of M87*, are illustrated in Figure 2.
@@
It is important to note that the precise profile of any DM spike is subject to significant theoretical uncertainties. A robust constraint on this model would require independent knowledge of the DM distribution. However, the primary goal of this work is to demonstrate a possible observational avenue. For a SMBH like M87* [EventHorizonTelescope:2019ggy] with a mass of $M\approx 6.5\times 10^{9}\mathrm{M}_{\odot}$ and $\gamma=1$, the DM density in the spike can be enhanced to extreme values, reaching up to $\rho_{\mathrm{max}}\sim\mathcal{O}(10^{12})\,\mathrm{GeV/cm^{3}}$ near the BH. This density can boost the DE effective mass to $\mu_{\mathrm{eff}}^{\mathrm{max}}\sim\mathcal{O}(10^{-24})\beta^{1/2}\,\mathrm{eV}$. For superradiance to be efficient for M87*, the boson mass must be $\mu\sim(G_{N}M)^{-1}\sim 10^{-20}\,\mathrm{eV}$. This implies that a relatively large coupling, $\beta\sim\mathcal{O}(10^{8})$, would be required to enhance the DE mass to the necessary level. If we take $\gamma=2$, then $\beta$ would need to be about $\mathcal{O}(10^{6})$.
@@
A significant mass enhancement could trigger a powerful superradiant instability, rapidly spinning down the central BH. In this work, we use M87* as an illustrative example, with a mass of $6500.0\pm 700.0\times 10^{6}\mathrm{M}_{\odot}$ and a spin parameter of $0.9^{+0.05}_{-0.05}$ (95% CL) [Tamburini:2019vrf]. Thus, the observed high spin for M87* constrains the coupling parameter $\beta$  [^6] Although M87* is thought to be rapidly spinning, its exact spin rate remains uncertain..
@@
5.3 Numerical method for spike-induced superradiance
@@
$\left[\square_{\mathrm{Kerr}}-\mu^{2}_{\mathrm{eff}}(r)\right]\Phi(t,r,\theta,\phi)=0,$ (5.6)
@@
$\Phi(t,r,\theta,\phi)=e^{-i\omega t}e^{im\phi}R(r)S(\theta).$ (5.7)
@@
Figure 3: The real part $M\omega_{R}$ and imaginary part $M\omega_{I}$ of the frequency for the fundamental superradiant mode ($\ell=m=1$) and the $\gamma=1$ case. The frequencies are plotted as a function of the coupling parameter $\beta$ for three representative values of the BH spin: $\tilde{a}=0.95$ (top), $\tilde{a}=0.9$ (middle), and $\tilde{a}=0.85$ (bottom). A superradiant instability exists in the region where $\omega_{I}>0$.
To solve this system, we employ a hybrid numerical approach. The angular equation is standard and can be solved efficiently using the continued fraction method [Leaver:1985ax, Dolan:2007mj, Liu:2022ygf] to determine the eigenvalue $\Lambda_{\ell m}$ for a given $a\omega$. However, this method does not apply to the radial equation due to the non-trivial radial dependence of the effective mass term $\mu_{\mathrm{eff}}^{2}(r)$. We therefore solve the radial equation using a direct integration or “shooting” method [Molina:2010fb]. This involves numerically integrating the radial equation inwards from spatial infinity and outwards from the event horizon to a common matching point, imposing the appropriate boundary conditions for a quasi-bound state (purely outgoing waves at infinity and purely ingoing waves at the horizon). The details of the numerical method can be found in Appendix D. We then search for the complex frequency $\omega$ that allows for a smooth matching of the solution and its derivative at the matching point. For this analysis, we focus on the fastest-growing fundamental mode, $\ell=m=1$.
@@
The physical mechanism at play is analogous to the classic “BH bomb” scenario, where a massive field is trapped by a reflecting mirror placed around a rotating BH [cardosoMatterKerrBlack2013]. In our model, the dense DM spike creates a gravitational potential well that acts as the “mirror”, trapping the DE field modes. The coupling strength $\beta$ controls the “reflectivity” of this mirror. A critical value of $\beta$ is required to create a potential barrier high enough to effectively trap the modes and trigger the instability. As $\beta$ increases further, the trapping becomes more efficient and the instability growth rate $\omega_{I}$ increases. However, a larger $\beta$ also increases the effective mass and thus the real frequency $\omega_{R}$. The instability is quenched once $\omega_{R}$ becomes too large and violates the superradiance condition $\omega_{R}<m\Omega_{H}$.
5.4 Applying superradiance constraints
We now apply the statistical framework developed in Section 3 to constrain the DE-DM coupling parameter, $\beta$, using observations of the SMBH M87*. We adopt the mass and spin values for M87* from recent observations, which indicate a high spin of $\tilde{a}=0.9\pm 0.05$ [Tamburini:2019vrf]. Although measurements of M87* spin have not yet reached a full consensus, its consistently high value makes it an excellent candidate for superradiance constraints.
The methodology is analogous to the one previously described, but it is adapted for the specifics of the current model. A key distinction is that our parameter space is now one-dimensional, consisting solely of the coupling constant $\beta$, as the mass of the superradiating boson (the DE field) is primarily determined by $\beta$ and the local DM density.
@@ 5.4 Applying superradiance constraints
    We then check if this growth rate is fast enough to spin down the BH within its lifetime. We compare the superradiance timescale, $\tau_{\mathrm{SR}}=1/\Gamma_{\mathrm{SR}}$, against the relevant astrophysical timescale for M87*, which we take to be $\tau_{\mathrm{BH}}\approx 10^{9}$ years following [Davoudiasl:2019nlo].
@@
Applying the statistical framework described above to our mock M87* BH sample, we derive constraints on the DE-DM coupling parameter, $\beta$, for two distinct scenarios corresponding to the initial halo profile index: $\gamma=1$ (NFW) and a steeper profile with $\gamma=2$. We analyze these two cases separately, assuming a log-uniform prior for $\beta$ within the ranges of $10^{7}-10^{11}$ for the $\gamma=1$ case and $10^{5}-10^{9}$ for the $\gamma=2$ case.
Our main results are the posterior probability distributions for $\beta$, presented in Figure 4. The shape of the posterior effectively quantifies how each value of $\beta$ is constrained by the observation of a high spin for M87*. For values of $\beta$ below a critical threshold, $\beta_{c}$ (around $10^{5.4}$ for $\gamma=1$ and $10^{7.4}$ for $\gamma=2$ respectively), the model is consistent with the data, as the coupling is too weak to induce a significant superradiant instability for the majority of the BH posterior samples. As $\beta$ approaches this critical value, the superradiance rate becomes efficient for most of the high-spin samples, causing the posterior probability to drop sharply.
@@ 6 Conclusions and discussions
Acknowledgments
We thank Zhuan Ning, Zi-Yan Yuwen and Cheng-Jun Fang for insightful discussions and comments. This work is supported by the National Key Research and Development Program of China Grant No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12588101, No. 12235019, and No. 12447101, and the China Manned Space Program Grant No. CMS-CSST-2025-A01.
@@ Appendix A Derivation of the effective energy transfer rate $Q$ from field theory
In this Appendix, we provide the detailed derivation of the continuity equations and the effective energy transfer rate $Q$ from the field-theoretic Lagrangian presented in Section 2.2 and also in Section 5.
@@
$T_{\mu\nu}=T_{\mu\nu}^{(\phi)}+T_{\mu\nu}^{(\chi)}+T_{\mu\nu}^{\mathrm{M}}\,.$ (A.1)
@@
$T_{\mu\nu}^{(\phi)}$ $=\partial_{\mu}\phi\partial_{\nu}\phi-g_{\mu\nu}\left[\frac{1}{2}\partial^{\alpha}\phi\partial_{\alpha}\phi+V\left(\phi\right)+xW(\phi,\chi)\right],$ (A.2)
$T_{\mu\nu}^{(\chi)}$ $=\partial_{\mu}\chi\partial_{\nu}\chi-g_{\mu\nu}\left[\frac{1}{2}\partial^{\alpha}\chi\partial_{\alpha}\chi+U\left(\chi\right)+(1-x)W(\phi,\chi)\right]\,$, (A.3)
where $x\in[0,1]$ is a real number chosen so that the sum of the energy densities of $\phi$ and $\chi$ equals the total dark sector energy density. The energy densities and pressures for the fields in a homogeneous and isotropic Universe are then read from the $00$ and $ii$ components of these tensors, respectively,
$\rho_{\phi}$ $=\frac{1}{2}\dot{\phi}^{2}+V(\phi)+xW(\phi,\chi)$ $P_{\phi}$ $=\frac{1}{2}\dot{\phi}^{2}-V(\phi)-xW(\phi,\chi)$ (A.4)
$\rho_{\chi}$ $=\frac{1}{2}\dot{\chi}^{2}+U(\chi)+(1-x)W(\phi,\chi)$ $P_{\chi}$ $=\frac{1}{2}\dot{\chi}^{2}-U(\chi)-(1-x)W(\phi,\chi)\,.$ (A.5)
The remaining contribution from the Standard Model particles can be modeled by a ideal fulid with the energy density $\rho_{\mathrm{M}}$, pressure $P_{\mathrm{M}}$ and the four velocity $u_{\mu}$ in a form of $T_{\mu\nu}^{M}=\rho_{M}u_{\mu}u_{\nu}+P_{M}\left(u_{\mu}u_{\nu}+g_{\mu\nu}\right)$. With these definitions, the total energy density of the dark sector is simply $\rho_{D}=\rho_{\phi}+\rho_{\chi}$. The $00$ component of the Einstein equation gives us the Friedmann equation relating the Hubble parameter $H=\dot{a}/a$ so that
$3H^{2}=8\pi G_{N}(\rho_{D}+\rho_{\mathrm{M}})\,.$ (A.6)
@@
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}+W_{,\phi}$ $=0\,$, (A.7)
$\ddot{\chi}+3H\dot{\chi}+U_{,\chi}+W_{,\chi}$ $=0\,$, (A.8)
where a subscript ,i,i denotes a partial derivative with respect to the field $i$. Taking the time derivative of the energy densities and using the equations of motion, we can express them in the form of fluid continuity equations,
$\dot{\rho}_{\phi}+3H\left(\rho_{\phi}+P_{\phi}\right)$ $=\dot{\phi}(\ddot{\phi}+3H\dot{\phi}+V_{,\phi}\dot{\phi}+W_{,\phi}\dot{\phi})-(1-x)W_{,\phi}\dot{\phi}+xW_{,\chi}\dot{\chi}$ (A.9)
$=-(1-x)W_{,\phi}\dot{\phi}+xW_{,\chi}\dot{\chi}\equiv Q_{\phi},$ (A.10)
$\dot{\rho}_{\chi}+3H\left(\rho_{\chi}+P_{\chi}\right)$ $=\dot{\chi}(\ddot{\chi}+3H\dot{\chi}+U_{,\chi}\dot{\chi}+W_{,\chi}\dot{\chi})-xW_{,\chi}\dot{\chi}+(1-x)W_{,\phi}\dot{\phi}$ (A.11)
$=xW_{,\chi}\dot{\chi}+(1-x)W_{,\phi}\dot{\phi}\equiv Q_{\chi}\,.$ (A.12)
This derivation shows that with this particular choice for splitting the stress-energy tensor, we recover $Q_{\chi}=-Q_{\phi}$, consistent with the phenomenological model in Eq. 2.2. When we set $x=0$, we recover the result in the main text.
However, this relation is not general. If we were to adopt a different convention, the form of $Q$ would change. For example, if we instead defined [Aboubrahim:2024spa]
$T_{\mu\nu}^{(\phi)}$ $=\partial_{\mu}\phi\partial_{\nu}\phi-g_{\mu\nu}\left[\frac{1}{2}\partial^{\alpha}\phi\partial_{\alpha}\phi+V\left(\phi\right)+W(\phi,\chi)\right],$ (A.13)
$T_{\mu\nu}^{(\chi)}$ $=\partial_{\mu}\chi\partial_{\nu}\chi-g_{\mu\nu}\left[\frac{1}{2}\partial^{\alpha}\chi\partial_{\alpha}\chi+V\left(\chi\right)+W(\phi,\chi)\right]\,$, (A.14)

### v2

@@ V.2 DE superradiance induced by DM spike
$\Omega(\Phi)=1+\frac{1}{2}\beta\left(\frac{\Phi}{M_{\mathrm{Pl}}}\right)^{2}.$ (34)
@@
$\square_{\mathrm{Kerr}}\Phi=\mu^{2}_{\mathrm{eff}}(\mathbf{r})\Phi,$ (35)
@@
$\mu_{\mathrm{eff}}^{2}(\mathbf{r})=\mu_{0}^{2}+\frac{\beta}{M_{\mathrm{Pl}}^{2}}\rho_{\mathrm{DM}}(\mathbf{r})\,.$ (36)
@@
However, the immense gravitational pull of an SMBH can dramatically increase the DM density in its vicinity, forming a structure known as a DM spike, where $\rho_{\mathrm{spike}}\gg\rho_{\mathrm{local}}$ [61]. This extreme density enhancement provides a mechanism to significantly boost the DE effective mass. A DM halo with an initial central density profile $\rho(r)=\rho_{0}(r/r_{0})^{-\gamma}$ will form a spike around an SMBH with a profile given by [61]
$\rho_{\mathrm{sp}}(r)=\rho_{R}\Theta(r-4R_{s})\left(1-\frac{4R_{\mathrm{s}}}{r}\right)^{3}\left(\frac{R_{\mathrm{sp}}}{r}\right)^{\gamma_{\mathrm{sp}}},$ (37)
@@
where $R_{\mathrm{s}}=2GM$ is the Schwarzschild radius of the SMBH; $\gamma_{\mathrm{sp}}=(9-2\gamma)/(4-\gamma)$ which yields $2.25<\gamma_{\mathrm{sp}}<2.5$ for $0<\gamma<2$. $\rho_{R}$ and $R_{\mathrm{sp}}$ is the spike parameter which depend on $\gamma$ and the BH mass $M$ through the relation $R_{\mathrm{sp}}\left(\gamma,M\right)=\alpha_{\gamma}r_{0}\left(M/\left(\rho_{0}r_{0}^{3}\right)\right)^{1/(3-\gamma)}$ and $\rho_{R}=\rho_{0}\left(R_{\mathrm{sp}}/r_{0}\right)^{-\gamma}$, where the normalization $\alpha_{\gamma}$ is numerically derived for each power-law index $\gamma$, e.g., $\alpha_{1}\approx 0.1$ and $\alpha_{2}\approx 0.02$ [61]. We determine the initial $\rho_{0}$ and $r_{0}$ through the mass-velocity-dispersion relationship following [94] (see Appendix C for more details).
@@
Figure 2: Dark matter density profiles around a SMBH with the mass of M87* ($M=6.5\times 10^{9}\mathrm{M}_{\odot}$). The solid line represents the DM spike profile resulting from an initial NFW halo ($\gamma=1$, $\gamma_{\mathrm{sp}}=7/3$), while the dashed line shows the steeper spike resulting from an initial $\gamma=2$ halo ($\gamma_{\mathrm{sp}}=5/2$). The density is shown as a function of the radial coordinate $r$, covering the range from the event horizon $r_{+}$ to the spike radius $R_{\mathrm{sp}}$.
@@
In this work, we will consider two representative cases for the initial halo profile. The first is the standard Navarro-Frenk-White (NFW) profile [51], corresponding to $\gamma=1$, which results in a spike density that scales as $\rho_{\mathrm{sp}}(r)\propto r^{-7/3}$. As a second, more optimistic case for generating a strong superradiance signal, we consider a steeper initial profile with $\gamma=2$, leading to an even denser spike with $\rho_{\mathrm{sp}}(r)\propto r^{-5/2}$. Since a steeper initial profile produces a denser DM spike, a smaller value of the coupling $\beta$ is required to enhance the DE effective mass to the superradiance threshold. We will derive constraints for both the $\gamma=1$ and $\gamma=2$ scenarios. The resulting DM spike density profiles for these two cases, calculated using the mass of M87*, are illustrated in Figure 2.
@@
It is important to note that the precise profile of any DM spike is subject to significant theoretical uncertainties. A robust constraint on this model would require independent knowledge of the DM distribution. However, the primary goal of this work is to demonstrate a possible observational avenue. For a SMBH like M87* [6] with a mass of $M\approx 6.5\times 10^{9}\mathrm{M}_{\odot}$ and $\gamma=1$, the DM density in the spike can be enhanced to extreme values, reaching up to $\rho_{\mathrm{max}}\sim\mathcal{O}(10^{12})\,\mathrm{GeV/cm^{3}}$ near the BH. This density can boost the DE effective mass to $\mu_{\mathrm{eff}}^{\mathrm{max}}\sim\mathcal{O}(10^{-24})\beta^{1/2}\,\mathrm{eV}$. For superradiance to be efficient for M87*, the boson mass must be $\mu\sim(G_{N}M)^{-1}\sim 10^{-20}\,\mathrm{eV}$. This implies that a relatively large coupling, $\beta\sim\mathcal{O}(10^{8})$, would be required to enhance the DE mass to the necessary level. If we take $\gamma=2$, then $\beta$ would need to be about $\mathcal{O}(10^{6})$.
@@
A significant mass enhancement could trigger a powerful superradiant instability, rapidly spinning down the central BH. In this work, we use M87* as an illustrative example, with a mass of $6500.0\pm 700.0\times 10^{6}\mathrm{M}_{\odot}$ and a spin parameter of $0.9^{+0.05}_{-0.05}$ (95% CL) [112]. Thus, the observed high spin for M87* constrains the coupling parameter $\beta$  [^7] Although M87* is thought to be rapidly spinning, its exact spin rate remains uncertain..
@@
V.3 Numerical method for spike-induced superradiance
@@
$\left[\square_{\mathrm{Kerr}}-\mu^{2}_{\mathrm{eff}}(r)\right]\Phi(t,r,\theta,\phi)=0,$ (38)
@@
$\Phi(t,r,\theta,\phi)=e^{-i\omega t}e^{im\phi}R(r)S(\theta).$ (39)
@@
To solve this system, we employ a hybrid numerical approach. The angular equation is standard and can be solved efficiently using the continued fraction method [77, 50, 84] to determine the eigenvalue $\Lambda_{\ell m}$ for a given $a\omega$. However, this method does not apply to the radial equation due to the non-trivial radial dependence of the effective mass term $\mu_{\mathrm{eff}}^{2}(r)$. We therefore solve the radial equation using a direct integration or “shooting” method [91]. This involves numerically integrating the radial equation inwards from spatial infinity and outwards from the event horizon to a common matching point, imposing the appropriate boundary conditions for a quasi-bound state (purely outgoing waves at infinity and purely ingoing waves at the horizon). The details of the numerical method can be found in Appendix D. We then search for the complex frequency $\omega$ that allows for a smooth matching of the solution and its derivative at the matching point. For this analysis, we focus on the fastest-growing fundamental mode, $\ell=m=1$.
Figure 3: The real part $M\omega_{R}$ and imaginary part $M\omega_{I}$ of the frequency for the fundamental superradiant mode ($\ell=m=1$) and the $\gamma=1$ case. The frequencies are plotted as a function of the coupling parameter $\beta$ for three representative values of the BH spin: $\tilde{a}=0.95$ (top), $\tilde{a}=0.9$ (middle), and $\tilde{a}=0.85$ (bottom). The blue dashed line marks the superradiance threshold $\omega_{R}=m\Omega_{H}$. A superradiant instability exists in the region where $\omega_{I}>0$.
@@
The physical mechanism at play is analogous to the classic “BH bomb” scenario, where a massive field is trapped by a reflecting mirror placed around a rotating BH [39]. In our model, the dense DM spike creates a gravitational potential well that acts as the “mirror”, trapping the DE field modes [^8] While a non-vanishing interior density (e.g., a “corona” configuration) can quench superradiant instabilities in the plasma-driven models [49, 123], this does not apply here. Standard cold dark matter is strictly collisionless; particles crossing the capture radius inevitably fall into the black hole without collisional replenishment. This phase-space depletion guarantees a strictly empty cavity, preserving a robust instability.. The coupling strength $\beta$ controls the “reflectivity” of this mirror. A critical value of $\beta$ is required to create a potential barrier high enough to effectively trap the modes and trigger the instability. As $\beta$ increases further, the trapping becomes more efficient and the instability growth rate $\omega_{I}$ increases. However, a larger $\beta$ also increases the effective mass and thus the real frequency $\omega_{R}$. The instability is quenched once $\omega_{R}$ becomes too large and violates the superradiance condition $\omega_{R}<m\Omega_{H}$.
V.4 Applying superradiance constraints
We now apply the statistical framework developed in Section III to constrain the DE-DM coupling parameter, $\beta$, using observations of the SMBH M87*. We adopt the mass and spin values for M87* from recent observations, which indicate a high spin of $0.9^{+0.05}_{-0.05}$ (95% CL) [112]. The methodology is analogous to the one previously described, but it is adapted for the specifics of the current model. A key distinction is that our parameter space is now one-dimensional, consisting solely of the coupling constant $\beta$, as the mass of the superradiating boson (the DE field) is primarily determined by $\beta$ and the local DM density.
@@ V.4 Applying superradiance constraints
    We then check if this growth rate is fast enough to spin down the BH within its lifetime. We compare the superradiance timescale, $\tau_{\mathrm{SR}}=1/\Gamma_{\mathrm{SR}}$, against the relevant astrophysical timescale for M87*, which we take to be $\tau_{\mathrm{BH}}\approx 10^{9}$ years following [46].
@@
Applying the statistical framework described above to our mock M87* BH sample, we derive constraints on the DE-DM coupling parameter, $\beta$, for two distinct scenarios corresponding to the initial halo profile index: $\gamma=1$ (NFW) and a steeper profile with $\gamma=2$. For this analysis, we generated 1000 mock posterior samples for each scenario. We analyze these two cases separately, assuming a log-uniform prior for $\beta$ within the ranges of $10^{7}-10^{11}$ for the $\gamma=1$ case and $10^{5}-10^{9}$ for the $\gamma=2$ case.
Our main results are the posterior probability distributions for $\beta$, presented in Figure 4. The shape of the posterior effectively quantifies how each value of $\beta$ is constrained by the observation of a high spin for M87*. For values of $\beta$ below a critical threshold, $\beta_{c}$ (around $10^{7.4}$ for $\gamma=1$ and $10^{5.4}$ for $\gamma=2$ respectively), the model is consistent with the data, as the coupling is too weak to induce a significant superradiant instability for the majority of the BH posterior samples. As $\beta$ approaches this critical value, the superradiance rate becomes efficient for most of the high-spin samples, causing the posterior probability to drop sharply.
@@ VI Conclusions and discussions
For this second model, it is important to discuss the potential impact of non-linear backreaction on the environmental DM. As the DE field grows, it exerts a gradient force that could perturb the DM density profile. Such backreaction effects are notoriously severe in the analogous context of standard plasma-driven superradiance, where non-linearities rapidly quench the instability through the phenomenon of relativistic transparency—as the plasma is accelerated to relativistic velocities, the plasma frequency (the effective mass of photons) is suppressed by the Lorentz factor $\gamma$ [40, 24, 35]. Remarkably, our scalar-tensor model is inherently protected from this most detrimental quenching mechanism [83]. The superradiant scalar field acquires its effective mass from the trace of the DM energy-momentum tensor, $T\simeq-\rho_{0}$. Because the rest-mass density $\rho_{0}$ is a Lorentz scalar, the effective mass remains strictly unsuppressed by the $\gamma$ factor, thereby entirely evading relativistic transparency. Combined with the collisionless nature of DM, which precludes the collective hydrodynamic blow-out typical of collisional plasmas, the instability in our scenario is theoretically substantially more robust. Although determining the full non-linear dynamics requires numerical simulations, the distinct physical nature of this mechanism justifies our current static estimation as a meaningful proof of principle.
@@
Acknowledgements.
We thank Zhuan Ning, Zi-Yan Yuwen, and Cheng-Jun Fang for insightful discussions and comments. This work is supported by the National Key Research and Development Program of China Grant No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No.12547110, No. 12588101, No. 12235019, and No. 12447101, and the China Manned Space Program Grant No. CMS-CSST-2025-A01.
DATA AVAILABILITY
The data that support the findings of this article are openly available [90].
@@ Appendix A Derivation of the effective energy transfer rate $Q$ from field theory
In this Appendix, we provide the detailed derivation of the continuity equations and the effective energy transfer rate $Q$ from the field-theoretic Lagrangian presented in Section II.2 and also in Section V.
@@
$\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a^{2}(t)\,\mathrm{d}\mathbf{x}^{2},$ (42)
where $a$ is the time-dependent scale factor. An essential ingredient for studying cosmic evolution is the Einstein field equations,
$G_{\mu\nu}=8\pi G_{N}\,T_{\mu\nu}\,$, (43)
where the stress-energy tensor $T_{\mu\nu}$ contains contributions from the $\phi$ and $\chi$ dark sectors as well as the Standard Model particles,
$T_{\mu\nu}=T_{\mu\nu}^{(\phi)}+T_{\mu\nu}^{(\chi)}+T_{\mu\nu}^{\mathrm{M}}\,.$ (44)
@@
$T_{\mu\nu}^{(\phi)}$ $=\partial_{\mu}\phi\partial_{\nu}\phi-g_{\mu\nu}\left[\frac{1}{2}\partial^{\alpha}\phi\partial_{\alpha}\phi+V\left(\phi\right)+xW(\phi,\chi)\right],$ (45)
$T_{\mu\nu}^{(\chi)}$ $=\partial_{\mu}\chi\partial_{\nu}\chi-g_{\mu\nu}\left[\frac{1}{2}\partial^{\alpha}\chi\partial_{\alpha}\chi+U\left(\chi\right)+(1-x)W(\phi,\chi)\right]\,$, (46)
where $x\in[0,1]$ is a real number chosen so that the sum of the energy densities of $\phi$ and $\chi$ equals the total dark sector energy density. The energy densities and pressures for the fields in a homogeneous and isotropic Universe are then read from the 0000 and iiii components of these tensors, respectively,
$\rho_{\phi}$ $=\frac{1}{2}\dot{\phi}^{2}+V(\phi)+xW(\phi,\chi)\,$,
$P_{\phi}$ $=\frac{1}{2}\dot{\phi}^{2}-V(\phi)-xW(\phi,\chi)$
$\rho_{\chi}$ $=\frac{1}{2}\dot{\chi}^{2}+U(\chi)+(1-x)W(\phi,\chi)\,$,
$P_{\chi}$ $=\frac{1}{2}\dot{\chi}^{2}-U(\chi)-(1-x)W(\phi,\chi)\,.$ (47)
The remaining contribution from the Standard Model particles can be modeled by a ideal fulid with the energy density $\rho_{\mathrm{M}}$, pressure $P_{\mathrm{M}}$ and the four velocity $u_{\mu}$ in a form of $T_{\mu\nu}^{M}=\rho_{M}u_{\mu}u_{\nu}+P_{M}\left(u_{\mu}u_{\nu}+g_{\mu\nu}\right)$. With these definitions, the total energy density of the dark sector is simply $\rho_{D}=\rho_{\phi}+\rho_{\chi}$. The 0000 component of the Einstein equation gives us the Friedmann equation relating the Hubble parameter $H=\dot{a}/a$ so that
$3H^{2}=8\pi G_{N}(\rho_{D}+\rho_{\mathrm{M}})\,.$ (48)
@@
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}+W_{,\phi}$ $=0\,$, (49)
$\ddot{\chi}+3H\dot{\chi}+U_{,\chi}+W_{,\chi}$ $=0\,$, (50)
where a subscript ,i,i denotes a partial derivative with respect to the field ii. Taking the time derivative of the energy densities and using the equations of motion, we can express them in the form of fluid continuity equations,
$\dot{\rho}_{\phi}+3H\left(\rho_{\phi}+P_{\phi}\right)$
$=\dot{\phi}\bigl(\ddot{\phi}+3H\dot{\phi}+V_{,\phi}\dot{\phi}+W_{,\phi}\dot{\phi}\bigr)-(1-x)W_{,\phi}\dot{\phi}+xW_{,\chi}\dot{\chi}$
$=-(1-x)W_{,\phi}\dot{\phi}+xW_{,\chi}\dot{\chi}\equiv Q_{\phi},$
$\dot{\rho}_{\chi}+3H\left(\rho_{\chi}+P_{\chi}\right)$
$=\dot{\chi}\bigl(\ddot{\chi}+3H\dot{\chi}+U_{,\chi}\dot{\chi}+W_{,\chi}\dot{\chi}\bigr)-xW_{,\chi}\dot{\chi}+(1-x)W_{,\phi}\dot{\phi}$
$=xW_{,\chi}\dot{\chi}+(1-x)W_{,\phi}\dot{\phi}\equiv Q_{\chi}\,.$ (51)
This derivation shows that with this particular choice for splitting the stress-energy tensor, we recover $Q_{\chi}=-Q_{\phi}$, consistent with the phenomenological model in Eq. 2. When we set $x=0$, we recover the result in the main text.
However, this relation is not general. If we were to adopt a different convention, the form of $Q$ would change. For example, if we instead defined [2]
$T_{\mu\nu}^{(\phi)}$ $=\partial_{\mu}\phi\,\partial_{\nu}\phi-g_{\mu\nu}\left[\frac{1}{2}\partial^{\alpha}\phi\,\partial_{\alpha}\phi+V\left(\phi\right)+W(\phi,\chi)\right],$ (52)
$T_{\mu\nu}^{(\chi)}$ $=\partial_{\mu}\chi\,\partial_{\nu}\chi-g_{\mu\nu}\left[\frac{1}{2}\partial^{\alpha}\chi\,\partial_{\alpha}\chi+V\left(\chi\right)+W(\phi,\chi)\right]\,$, (53)